Shift report 20010820 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 20 2001 - 23:46:51 EDT

  • Next message: Flemming Videbaek: "SEB setup, RHIC maintenance on Thu Aug 23"

    ------------------------------------------------------------
    Date: Mon Aug 20 23:46:51 2001, Shift: 16:00-24:00
    Supervisor: JJG, with: AM
    Log book#7 pages 144 - 145
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam during entire shift. 22-23PM:two magnet quenches.
       RHIC is recovering. Hope the shifters will too.
    2. At start of shift, after the storm and the flood, D1 
       and D2 could not be powered. Called MCD. After 45mn of
       various hard and soft resets all magnets are operational.
    3. All HV (except know troublemakers) are operational. 
       DC are currently off and TPC's down to stby volt.
    4. Achievements: none, except for posting a list of the DC 
       voltages on the CH wall (inner) and some general cleaning.
    5. Various suggestions:
       a) If a single item on this shift report is omitted, 
          the entire text wanishes, and must be retyped fromn 
          scratch!!!
       b) It takes about 20-25 mn to raise the TPC and DC voltages
          at beginning of a run. This is a major fraction of the
          beam time. Will someone implement the general BRAHVO buttons 
          to lower and raise the HV on DC+TPC in the fastest 
          possible way?
       c) I have noted that our presentations are not public
          on the BRAHMS webpage. Suggest to move that section
          to the public part ASAP!
    6. Summary: a boring evening.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 20 2001 - 23:47:28 EDT